Ann can solve 48 problems in 21 minutes, and Sara can solve 36 problems in 84 minutes . How many problems can Ann Solve in the same amount of time it takes Sara to solve 54 problems.
A)72
B)144
C)192
C)288

Answer by rothauserc(4718)

36/84 = 54/x
:
36x = 84 * 54 = 4536
:
x = 126 minutes
:
48/21 = y/126
:
21y = 48 * 126 = 6048
:
y = 288
:
****************
answer is C) 288
****************
: